# Optical Physiological Signals

### v1

@@ Abstract
Photoplethysmography (PPG) is the most widely used non-invasive technique for monitoring biosignals and cardiovascular health, with applications in both clinical settings and consumer health through wearable devices. Current machine learning models trained on PPG signals are mostly task-specific and lack generalizability. Previous works often used single-device datasets, did not explore out-of-domain generalization, or did not release their models, hindering reproducibility and further research. We introduce PaPaGei, the first open foundation model for PPG signals. PaPaGei is pre-trained on more than 57,000 hours of 20 million unlabeled segments of PPG signals using publicly available datasets exclusively. We evaluate against popular time-series foundation models and other benchmarks on 20 tasks of 10 diverse datasets spanning cardiovascular health, sleep disorders, pregnancy monitoring, and wellbeing assessment. Our architecture incorporates novel representation learning approaches that leverage differences in PPG signal morphology across individuals, enabling it to capture richer representations than traditional contrastive learning methods. Across 20 tasks, PaPaGei improves classification and regression performance by an average of 6.3% and 2.9%, respectively, compared to other competitive time-series foundation models in at least 14 tasks. PaPaGei is more data- and parameter-efficient than other foundation models or methods, as it outperforms 70x larger models. Beyond accuracy, we also investigate robustness against different skin tones, establishing a benchmark for bias evaluations of future models. Notably, PaPaGei can be used out of the box as both a feature extractor and an encoder for other multimodal models, opening up new opportunities for multimodal health monitoring [^1] Models, data pipelines, and code are available at: https://github.com/nokia-bell-labs/papagei-foundation-model.
@@ 1 Introduction
Photoplethysmography (PPG), a non-invasive optical sensing technique, is widely used for monitoring cardiovascular health and physiological signals in both clinical and consumer health applications (Charlton et al., 2023). From hospital pulse oximeters to smartwatches, PPG enables continuous health monitoring in various settings, bridging acute medical care and long-term health management. PPG signals can track a diverse range of health indicators, including cardiovascular health, blood pressure, mood, and sleep disorders (Sadad et al., 2022; Ave et al., 2015; Temko, 2017; Reiss et al., 2019; Liang et al., 2018a; Haddad et al., 2021; Schrumpf et al., 2021).
Despite its widespread adoption, PPG presents significant challenges for machine learning (ML). These include high annotation costs due to the need for domain-specific expertise, especially in consumer health applications where recording conditions and subject heterogeneity vary greatly. PPG signals are also susceptible to noise and motion artifacts (Afandizadeh Zargari et al., 2023), and inherent variability due to factors like skin tone and body composition (Bent et al., 2020) complicates the development of generalizable ML models. Consequently, existing PPG datasets are often small, task-specific, and limited in their generalizability, posing a major obstacle to the development of robust and widely applicable models that could fully leverage the potential of this technology.
@@
The PPG domain lacks general-purpose foundation models, unlike language and vision domains, with most work focused on single-dataset task-specific models. Although PPG signals detect vital signs like heart rate variability and blood oxygen saturation (Abbaspourazad et al., 2023; Charlton et al., 2023), the absence of generalizable pre-trained models limits progress. While there are difficulties in obtaining large-scale, high-quality datasets, recent increases in heterogeneous PPG data offer new opportunities (Johnson et al., 2016; Zhang et al., 2018; Lee et al., 2022). To address these challenges, we introduce PaPaGei, a set of robust, pre-trained models capable of serving as a backbone for various PPG-related tasks, capturing rich PPG representations through large-scale pre-training.
Our key contributions are: (1) Large-scale pre-training for PPG signals: To our knowledge, PaPaGei is the first open foundation model pre-trained on PPG signals, using 57,000 hours of data from 20 million signals sourced entirely from public datasets. This establishes a new benchmark for large-scale model development in wearable and clinical health monitoring. (2) PPG-aware self-supervised learning framework: We introduce a novel self-supervised learning framework with a unique PPG signal morphology augmentation module. Our approach optimizes agreement between PPG signals with similar blood volume changes while learning signal quality and ratio around the dicrotic notch, a key PPG marker. (3) Comprehensive evaluation across diverse out-of-domain health tasks: We evaluate PaPaGei across 20 tasks, including cardiovascular health, sleep disorders, pregnancy monitoring, and overall well-being. Our results show that the model’s embeddings contain rich and predictive information applicable to various health conditions, outperforming existing benchmarks. (4) Extensive robustness studies: We conduct ablation studies to assess the impact of key components, including signal morphology augmentation, comparisons with established contrastive learning approaches, model size, data efficiency, and the effect of skin tone.
@@ 2 Related work
Self-supervised learning (SSL) has become a prominent paradigm for learning general representations from unlabeled datasets, with applications in physiological signal analysis including health, fitness, and brain signals (Tonekaboni et al., 2021; Zhang et al., 2022; Chen et al., 2021; Yèche et al., 2021; Spathis et al., 2021; Cheng et al., 2020; Kiyasseh et al., 2021; Sarkar & Etemad, 2020). Despite SSL’s popularity, there are no widely used pre-trained models for PPG signals. Recent studies have shown promise in this area, as (Abbaspourazad et al., 2023) demonstrated that embeddings from ECG and PPG signals can generalize across multiple health-related tasks using proprietary Apple Watch data. Similarly, (Yun et al., 2024) showed that embedding PPG signals can improve genetic discovery and risk prediction outcomes using the UK Biobank dataset. Other works (Weng et al., 2024; Ding et al., 2024; Zhou et al., 2024) explored PPG embeddings for various applications. However, these studies often used single-device datasets, did not explore out-of-domain generalization, or did not release their models, highlighting the need for openly available, pre-trained PPG foundation models.
While generic time-series foundation models like Chronos (Ansari et al., 2024) and Moment (Goswami et al., 2024) have gained popularity, they often lack significant physiological data representation. In contrast, our work takes a domain-specific approach, focusing exclusively on PPG signals. We argue that while knowledge from generic models may transfer to PPG tasks, performance is likely to be limited compared to a PPG-specific model. Similarly, transferring from other domain-specific models (e.g., ECG (McKeen et al., 2024; Song et al., 2024) or EEG (Yuan et al., 2024b)) to PPG is challenging due to differences in signal characteristics. Our specialized approach aims to capture the nuances and complexities specific to PPG signals, building on the increasing interest in modality-specific foundation models. See Appendix §F for more detailed discussions.
Figure 2: PaPaGei-S Overview. Before training, for each PPG segment we compute the morphology metrics IPA, SVRI, and SQI, to be used as objectives (left). Next, we pass raw PPG signals through the encoder ($E$), to extract embeddings (middle) and feed them into three heads (right): the projection head ($P$) contrasts similar and dissimilar PPG signals based on sVRI, while two mixture-of-expert heads ($M_{1}$ and $M_{2}$) refine the embeddings by predicting IPA and SQI values, respectively.
@@ 3 Methods
Given a dataset $\mathcal{D}=\{\mathbf{p}^{1},\mathbf{p}^{2},\cdots,\mathbf{p}^{S}\}$ representing diverse PPG signals from $S$ subjects, a PPG signal $\mathbf{p}^{s}\in\mathbb{R}^{n}$ is defined as a time series that models the changes in light intensity due to arterial blood flow. To model granular changes in PPG signal of each subject $s$, we segment $\mathbf{p}^{s}$ without overlap to obtain $X^{s}=\{\mathbf{x}_{1}^{s},\mathbf{x}_{2}^{s},\cdots\mathbf{x}_{N}\}$. Here, the number of segments $N$ depends on the sampling frequency ($f$) and the desired length of time window.
We use self-supervised learning to train PaPaGei’s foundation model. We propose PaPaGei-P, a patient contrastive approach that maximizes agreement between signals from the same subject. Importantly, we propose PaPaGei-S, a morphology-aware self-supervised approach that maximizes agreement between signals with similar morphology. We present the details of both approaches.
3.1 Participant-aware objective
@@
In PaPaGei-P, we train a self-supervised learning model to maximize agreement between the latent embeddings of PPG signals from the same subject. While (Abbaspourazad et al., 2023) implements a similar strategy, they do not train or evaluate on public datasets.
@@
Training. We define a positive pair as any two distinct segments from the same subject, denoted as $\{(\mathbf{x}^{s}_{i},\mathbf{x}^{s}_{j})|i\neq j\}$. Next, we apply a series of time series augmentations such as random cropping, adding Gaussian noise, time flipping, negation, and magnitude scaling (Tang et al., 2020). Each augmentation is applied with a predefined probability, determining its likelihood during training. Additionally, each augmentation includes hyper-parameters that control the intensity of the transformation. During training, the augmented versions of two randomly sampled positive pairs are passed through the encoder $E$ and projection $P$ to obtain embeddings denoted $(\mathbf{z}_{i}^{s},\mathbf{z}_{j}^{s})$. Given a batch of embeddings from $N$ distinct subjects with positive pairs of the form $(\mathbf{z}_{i}^{s},\mathbf{z}_{j}^{s})$, the model optimizes the NT-Xent loss given by: $\mathcal{L}_{p}=\frac{1}{2}(\ell_{p}(i,j)+\ell_{p}(j,i))$, where $\ell_{p}(i,j)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(sim(\mathbf{z}_{i}^{s},\mathbf{z}_{j}^{s})/\tau)}{\sum_{s\neq r}\exp(sim(\mathbf{z}_{i}^{s},\mathbf{z}_{j}^{r})/\tau)}$. In contrast, vanilla SimCLR (Chen et al., 2020) uses positive pairs as augmented versions of randomly sampled PPG segments.
3.2 Morphology-aware objective
@@
In PaPaGei-S, we leverage the PPG signal morphology to train a self-supervised learning model that maximizes agreement between similar physiological features of PPG signals across subjects.
@@
PPG Morphology. Total peripheral resistance (TPR)—the force exerted by the body’s blood vessels on circulating blood—varies under certain medical conditions, such as hypertension and diabetes (Trammel & Sapra, 2020). Variations in TPR are reflected in PPG signals, presenting as distinct regions within the waveform (Figure 2, left). To capture these variations, we introduce a morphology augmentation module before training, which computes three key PPG metrics (Figure 2, left): (1) stress-induced Vascular Response Index (sVRI) (Lyu et al., 2015; Zhang et al., 2019): the ratio of mean PPG signal between post- to pre-systolic phases, (2) Inflection Point Area ratio (IPA) (Wang et al., 2009): the ratio of systolic to diastolic areas defined by the dicrotic notch, and (3) Signal Quality index (SQI): skewness of the signal as an indicator of quality (Elgendi, 2016). We choose these metrics because they complement each other, with sVRI capturing amplitude variations, IPA reflecting signal width, and SQI addressing cases where IPA cannot be computed due to poor-quality signals lacking a dicrotic notch. We compute these metrics as follows:
@@
where $\mathbf{x}\in\mathbb{R}^{N}$ is the PPG segment, $sys$ is the systolic peak, $n$ is the length of time series, and $\hat{n}$ is the dicrotic notch. For $SQI$, we divide x into 5 second windows ($w$; total windows $W$) and compute the skewness $m_{i}=\frac{1}{5\times f}\sum_{j=1}^{5\times f}(x[j]-\mu_{x}[j])^{i}$, which gives the best signal quality discrimination (Elgendi, 2016).
Training. Before training, the morphology augmentation module takes an augmented (only Gaussian noise and cropping) input time series $\mathbf{x}$ and outputs $y=\{y^{svri},y^{ipa},y^{sqi}\}\in\mathbb{R}^{3}$ (Figure 2 middle). Next, we discretize $y^{svri}$ into a predefined set of $n=8$ bins to denote positive pairs, where $y^{svri}\in\{0,1,\dots,n\}$. We define positive pairs based on the sVRI labels as $\{(\mathbf{x}_{i},\mathbf{x}_{j})|y^{svri}_{i}=y^{svri}_{j},i\neq j\}$. Note that positive pairs are not defined based on subjects.
Given a batch of $N$ PPG signals and their morphology, we optimize three heads. First, we extract the embeddings $Z=\{\textbf{z}_{1},\textbf{z}_{2},\cdots,\textbf{z}_{N}\}$ from the projection $P$ (Figure 2 right middle), and compute the contrastive loss for sVRI (equation 2). Next, we use the embeddings $H=\{\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{N}\}$ to predict the IPA ($\mathbf{\hat{y}}^{ipa}\in\mathbb{R}^{N}$) and SQI ($\mathbf{\hat{y}}^{sqi}\in\mathbb{R}^{N}$) using the mixture of expert (MoE) heads $M_{1}$ and $M_{2}$. Each MoE head is composed of three fully connected neural networks (FCNNs), with the head’s output calculated as a weighted sum of the FCNNs, using softmax to determine the weights. These heads are optimized using the mean absolute error (equation 3). The morphology indices encapsulate various PPG characteristics. Our rationale for utilizing MoE is that each expert can specialize in learning distinct properties that contribute to the overall index. Finally, the overall PaPaGei-S training objective is given in equation 4.
$\ell_{s}(i,j)$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(sim(\mathbf{z}_{i},\mathbf{z}_{j})/\tau\right)}{\sum_{k\neq j}\exp\left(sim(\mathbf{z}_{i},\mathbf{z}_{k})/\tau\right)}$ (2)
$\mathcal{L}_{svri}=\frac{1}{2}(l_{s}(i,j)+l_{s}(j,i))\quad$ $\mathcal{L}_{ipa}=\frac{1}{N}\sum_{i=1}^{N}\left|y^{ipa}_{i}-\hat{y}^{ipa}_{i}\right|\quad\mathcal{L}_{sqi}=\frac{1}{N}\sum_{i=1}^{N}\left|y^{sqi}_{i}-\hat{y}^{sqi}_{i}\right|$ (3)

### v2

@@ Abstract
Photoplethysmography (PPG) is the leading non-invasive technique for monitoring biosignals and cardiovascular health, with widespread adoption in both clinical settings and consumer wearable devices. While machine learning models trained on PPG signals have shown promise, they tend to be task-specific and struggle with generalization. Current research is limited by the use of single-device datasets, insufficient exploration of out-of-domain generalization, and a lack of publicly available models, which hampers reproducibility. To address these limitations, we present PaPaGei, the first open foundation model for PPG signals. The model is pre-trained on over 57,000 hours of data, comprising 20 million unlabeled PPG segments from publicly available datasets. We introduce a novel representation learning approach that leverages domain knowledge of PPG signal morphology across individuals, enabling the capture of richer representations compared to traditional contrastive learning methods. We evaluate PaPaGei against state-of-the-art time-series foundation models and self-supervised learning benchmarks across 20 tasks from 10 diverse datasets, spanning cardiovascular health, sleep disorders, pregnancy monitoring, and wellbeing assessment. Our model demonstrates superior performance, improving classification and regression metrics by 6.3% and 2.9% respectively in at least 14 tasks. Notably, PaPaGei achieves these results while being more data- and parameter-efficient, outperforming models that are 70$\times$ larger. Beyond accuracy, we examine model robustness across different skin tones, establishing a benchmark for bias evaluation in future models. PaPaGei can serve as both a feature extractor and an encoder for multimodal models, opening up new opportunities for multimodal health monitoring [^1] Models, data, and code are available at: github.com/nokia-bell-labs/papagei-foundation-model.
@@ 1 Introduction
Photoplethysmography (PPG), a non-invasive optical sensing technique, is widely used to monitor cardiovascular health and physiological signals in both clinical and consumer health applications (Charlton et al., 2023). From hospital pulse oximeters to smartwatches, PPG enables continuous health monitoring in various settings, bridging acute medical care and long-term health management. PPG signals help in tracking a diverse range of health indicators, including cardiovascular health, blood pressure, mood, and sleep disorders (Sadad et al., 2022; Ave et al., 2015; Reiss et al., 2019; Liang et al., 2018a; Haddad et al., 2021; Schrumpf et al., 2021). Despite its widespread adoption, PPG poses substantial challenges for machine learning applications. A primary obstacle is the high cost of data annotation, which requires specialized domain expertise. This challenge is particularly pronounced in consumer health applications, where varying sensing conditions and diverse user populations create additional complexity. PPG signals are susceptible to noise and motion artifacts (Afandizadeh Zargari et al., 2023), as well as inherent variability due to factors like skin tone and body composition (Bent et al., 2020). These complicate the development of generalizable ML models for PPG. Consequently, existing PPG datasets are often small, task-specific, and limited in their generalizability, posing a major obstacle to the development of robust and widely applicable models that could fully leverage the potential of PPG technology.
@@
The PPG domain, unlike language or vision domains, lacks general-purpose foundation models (FMs), with most current works focused on single-dataset task-specific models. Although PPG can detect vital signs like heart rate variability and blood oxygen saturation, the absence of generalizable pre-trained models limits progress (Abbaspourazad et al., 2023). Despite the ongoing challenges of acquiring large-scale, high-quality data, recent expansions in diverse PPG datasets have created new opportunities (Johnson et al., 2016; Zhang et al., 2018; Lee et al., 2022). To address these challenges, we introduce PaPaGei, a set of robust, pre-trained models capable of serving as a backbone for various PPG-related tasks, capturing rich PPG representations through large-scale pre-training.
The key contributions of PaPaGei are:
(1) Large-scale pre-training for PPG signals: To our knowledge, PaPaGei is the first open foundation model pre-trained on PPG signals, using 57,000 hours of data from 20 million signals sourced entirely from public datasets. This establishes a new benchmark for large-scale model development in wearable and clinical health monitoring.
(2) PPG-aware self-supervised learning (SSL) framework: We introduce a novel SSL framework with a unique PPG signal morphology augmentation module. Our approach optimizes agreement between PPG signals with similar blood volume changes while pointing the model to pay attention to the changes around the systolic peak and dicrotic notch (key PPG markers).
(3) Comprehensive evaluation across diverse out-of-domain health tasks: We evaluate PaPaGei across 20 tasks, including cardiovascular health, sleep disorders, pregnancy monitoring, and overall well-being. Our results show that the model embeddings contain rich and predictive information applicable to various health conditions, outperforming existing benchmarks.
(4) Extensive robustness studies: We conduct ablation studies to assess the impact of key components, including signal morphology augmentation, comparisons with established contrastive learning approaches, model size, data efficiency, and the effect of skin tone.
@@ 2 Related work
Self-supervised learning has become a prominent paradigm for learning general representations from unlabeled datasets, with applications in physiological signal analysis including health, fitness, and brain signals (Tonekaboni et al., 2021; Zhang et al., 2022; Chen et al., 2021; Yèche et al., 2021; Spathis et al., 2021; Cheng et al., 2020; Kiyasseh et al., 2021; Sarkar & Etemad, 2020). Despite its popularity, there are no widely used models for PPG signals pre-trained through SSL. Recently, Abbaspourazad et al. (2023) demonstrated that embeddings derived from PPG signals can predict over 45 diverse downstream health-related tasks using proprietary Apple Watch data. Their approach uses an SSL framework based on patient-level positive pair contrastive learning. Similarly, (Yun et al., 2024) showed that embedding PPG signals can improve genetic discovery and risk prediction outcomes using the UK Biobank dataset. Other works (Weng et al., 2024; Ding et al., 2024; Zhou et al., 2024) explored PPG embeddings for various applications. However, these studies often used proprietary datasets, did not explore out-of-domain generalization, or did not release their models, highlighting the need for openly available, pre-trained PPG FMs (Table 17). For example, in contrast to (Abbaspourazad et al., 2023), our work exclusively uses public datasets for large-scale PPG training and introduces a novel SSL framework to incorporate PPG morphology. While Abbaspourazad et al. (2023) evaluate a single proprietary dataset, we validate on 10 diverse downstream datasets, showcasing greater generalizability and robustness across varied real-world scenarios.
Generic time-series FMs, like Chronos (Ansari et al., 2024) and Moment (Goswami et al., 2024), lack physiological data representation. There is growing interest in modality-specific FMs tailored to physiological signals (Song et al., 2024; Lai et al., 2023) and human activity (Yuan et al., 2024a). Knowledge transfer from time-series FMs might benefit PPG tasks, but their performance is limited compared to PPG-specific FMs. Adapting other domain-specific models, like ECG (McKeen et al., 2024; Song et al., 2024) or EEG (Yuan et al., 2024b), is challenging due to distinct signal characteristics. We specifically design FMs for PPG signals, contributing to the growing movement toward foundation models tailored to individual modalities. See Appendix §H for an extended discussion.
@@ 3 Methods
Figure 2: Overview of PaPaGei-S. The process begins by computing three morphology metrics (IPA, SVRI, and SQI) for each PPG segment. The raw PPG signals are then processed through an encoder ($E$) to generate embeddings ($H$). These same embeddings feed into three specialized heads: a projection head ($P$) that contrasts PPG signals based on sVRI values, and two mixture-of-expert heads ($M_{1}$ and $M_{2}$) that refine the embeddings by predicting IPA and SQI values.
@@
Given a dataset $\mathcal{D}=\{\mathbf{p}^{1},\mathbf{p}^{2},\cdots,\mathbf{p}^{S}\}$ representing diverse PPG signals from $S$ participants, a PPG signal $\mathbf{p}^{s}\in\mathbb{R}^{n}$ is defined as a time-series that captures variations in light intensity caused by arterial blood flow. To model granular changes in PPG signal of each subject $s$, we segment $\mathbf{p}^{s}$ without overlap to obtain $X^{s}=\{\mathbf{x}_{1}^{s},\mathbf{x}_{2}^{s},\cdots\mathbf{x}^{s}_{N}\}$. Here, the number of segments $N$ depends on the sampling frequency ($f$) and the desired length of time window. To train our foundation models, PaPaGei-P employs a patient contrastive SSL approach that maximizes agreement between signals from the same subject. Importantly, we propose PaPaGei-S, a morphology-aware self-supervised approach that maximizes agreement between PPG segments with similar morphology.
3.1 Participant-aware objective: PaPaGei-P
@@
In PaPaGei-P, we train an SSL model to maximize agreement between the embeddings of PPG signals from the same subject. While previous studies have demonstrated the effectiveness of this strategy for physiological signals (Kiyasseh et al., 2021; Abbaspourazad et al., 2023), our work represents the first attempt to train and evaluate a foundation model using publicly available PPG datasets.
@@
Training. We define a positive pair as any two distinct segments of PPG signals from the same subject, denoted as $\{(\mathbf{x}^{s}_{i},\mathbf{x}^{s}_{j})|i\neq j\}$. Next, we apply a series of time-series augmentations such as random cropping, adding Gaussian noise, time flipping, negation, and magnitude scaling (Tang et al., 2020), each applied with a predefined probability during training. Each augmentation includes hyper-parameters that control the intensity of the data transformation. During training, the augmented version of a randomly sampled positive pair $(\mathbf{x}^{s}_{i},\mathbf{x}^{s}_{j})$ is passed through the encoder $E$, and subsequently projection $P$, to obtain an embeddings pair denoted by $(\mathbf{z}_{i}^{s},\mathbf{z}_{j}^{s})$. Given a batch of embeddings from $N$ positive pairs of the form $(\mathbf{z}_{i},\mathbf{z}_{j})$, the model optimizes the normalized temperature-scaled cross entropy (NT-Xent) loss (Sohn, 2016; Oord et al., 2018; Chen et al., 2020) given by: $\mathcal{L}_{p}=\frac{1}{2}(\ell_{p}(i,j)+\ell_{p}(j,i))$, where $\ell_{p}(i,j)=-\frac{1}{N}\sum_{u=1}^{N}\log\frac{\exp(sim(\mathbf{z}_{i}^{u},\mathbf{z}_{j}^{u})/\tau)}{\sum_{v=1}^{2N}\mathbbm{1}[v\neq u]\exp(sim(\mathbf{z}_{i}^{u},\mathbf{z}_{j}^{v})/\tau)}$ and $sim(\cdot,\cdot)$ is the cosine similarity. In contrast, vanilla SimCLR (Chen et al., 2020) would use positive pairs as augmented versions of randomly sampled PPG segments.
3.2 Morphology-aware objective: PaPaGei-S
@@
In PaPaGei-S, we leverage the PPG signal morphology to train a SSL model that maximizes agreement between similar physiological features of PPG signals across participants.
@@
PPG Morphology. Total peripheral resistance (TPR)—the force exerted by the body’s blood vessels on circulating blood—varies under certain medical conditions, such as hypertension and diabetes (Trammel & Sapra, 2020). Variations in TPR are reflected in PPG signals, presenting as distinct regions within the waveform. To capture these variations, we introduce a morphology augmentation module before training, which computes three key PPG metrics (Figure 2, left): (1) stress-induced Vascular Response Index (sVRI) (Lyu et al., 2015; Zhang et al., 2019): the ratio of mean PPG signal between post- to pre-systolic phases, (2) Inflection Point Area ratio (IPA) (Wang et al., 2009): the ratio of systolic to diastolic areas defined by the dicrotic notch, and (3) Signal Quality Index (SQI): skewness of the signal as an indicator of quality (Elgendi, 2016). Prior studies have shown that incorporating the PPG signal quality during training yields positive results (Ding et al., 2024). We selected these metrics for their complementary nature: sVRI captures variations in amplitude, while IPA measures signal width. To address scenarios where computing IPA is challenging because of noisy signals or different morphology, we incorporate SQI. In particular, we empirically find that SQI is significantly larger ($p<0.05$) in signals with a dicrotic notch (Appendix §D.5).
@@
where $\mathbf{x}\in\mathbb{R}^{N}$ is the PPG segment, $sys$ is the systolic peak, $n$ is the length of time series, and $\hat{n}$ is the dicrotic notch. For $SQI$, we divide x into 5 second windows ($w$; total windows $W$) and compute the skewness $m_{i}=\frac{1}{5\times f}\sum_{j=1}^{5\times f}(x[j]-\mu_{x}[j])^{i}$, which gives the best signal quality discrimination.
Training. Before training, the morphology augmentation module takes an augmented input, by applying Gaussian noise and cropping to time series $\mathbf{x}$, and outputs $y=\{y^{svri},y^{ipa},y^{sqi}\}\in\mathbb{R}^{3}$ (Figure 2 middle). Next, we discretize $y^{svri}$ into a predefined set of $b=8$ bins to denote positive pairs, where $y^{svri}\in\{1,\dots,b\}$. We define positive pairs based on the sVRI labels as $\{(\mathbf{x}_{i},\mathbf{x}_{j})|y^{svri}_{i}=y^{svri}_{j},i\neq j\}$. Note that positive pairs are not defined based on participants.
$\ell_{s}(i,j)=-\log\frac{\exp\left(sim(\mathbf{z}_{i},\mathbf{z}_{j})/\tau\right)}{\sum_{k=1}^{2N}\mathbbm{1}[k\neq i]\exp\left(sim(\mathbf{z}_{i},\mathbf{z}_{k})/\tau\right)}$ (2)
$\mathcal{L}_{svri}=\frac{1}{2N}\sum_{k=1}^{N}[\ell(2k-1,2k)+\ell(2k,2k-1)]\quad$ (3)
$\mathcal{L}_{ipa}=\frac{1}{N}\sum_{i=1}^{N}\left|y^{ipa}_{i}-\hat{y}^{ipa}_{i}\right|\quad\mathcal{L}_{sqi}=\frac{1}{N}\sum_{i=1}^{N}\left|y^{sqi}_{i}-\hat{y}^{sqi}_{i}\right|$ (4)
@@
Given a batch of $N$ PPG signals and their morphology, we optimize three heads using the encoder ($E$) embeddings $H=\{\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{N}\}$. First, we extract the embeddings $Z=\{\textbf{z}_{1},\textbf{z}_{2},\cdots,\textbf{z}_{N}\}$ from the projection ($P$), and compute the contrastive loss for sVRI (equation 3). Next, we use the embeddings $H$ to predict the IPA ($\mathbf{\hat{y}}^{ipa}\in\mathbb{R}^{N}$) and SQI ($\mathbf{\hat{y}}^{sqi}\in\mathbb{R}^{N}$) using the mixture of expert (MoE) heads $M_{1}$ and $M_{2}$. Each MoE head is composed of three fully connected neural networks (FCNNs), with the head’s output calculated as a weighted sum of the FCNNs, using softmax to determine the weights. These heads are optimized using the mean absolute error (equation 4). The morphology indices encapsulate various PPG characteristics. Our rationale for utilizing MoE is that each expert can specialize in learning distinct properties that contribute to the overall index. Finally, the overall PaPaGei-S training objective is given in equation 5.